OpenStudy (anonymous):

Let f(x)=x^2+5 and g(x) = x-7. Perform each operation and then find out the domain. 1) f(x)/g(x) 2) f(x)-2g(x) 3) f(x)*g(x)

OpenStudy (mertsj):

\[\frac{f(x)}{g(x)}=\frac{x^2+5}{x-7}\]

OpenStudy (mertsj):

Domain = all real numbers except 7

OpenStudy (mertsj):

\[f(x)-2g(x)=x^2+5-2(x-7)=x^2+5-2x-14=x^2-2x-9\]

OpenStudy (mertsj):

Domain = all real numbers

OpenStudy (mertsj):

\[f(x)\times g(x)=(x^2+5)(x-7)=x^3-7x^2+5x-35\]

OpenStudy (mertsj):

Domain = all real numbers